OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. x/x+3 + 2/5 = 0

OpenStudy (anonymous):

\[\frac{x}{x+3}+\frac{2}{5}=0\] \[\frac{x}{x+3}=-\frac{2}{5}\] \[5x=-2(x+3)\] and go from there